# Local converse theorems and Langlands parameters

### v4

@@ Abstract.
Let $F$ be a non Archimedean local field, and $G$ be the $F$-points of a connected quasi-split reductive group defined over $F$. In this paper we prove a converse theorem statement for generic Langlands parameters of $G$ when $G$ is $F$-split and the Langlands dual group of $G$ is acceptable, and propose a possible extension of it without the split assumption. We also prove that the statement does not apply to ${\mathrm{SO}}_{6}(F)$ for certain choices of $F$. Then we consider a variant which we prove for $G=\mathrm{G}_{2}(F)$ and all quasi-split classical groups. When $F$ has characteristic zero and assuming the validity of the Gross-Prasad and Rallis conjecture, this latter variant translates via the generic local Langlands correspondence of Jantzen and Liu, into the usual local converse theorems for classical groups expressed in terms of Shahidi’s gamma factors.
@@ 1. Introduction
We now briefly explain the main observations that are made in this paper. We denote by $G$ be the $F$-points of a connected quasi-split reductive group defined over $F$, with Weil-Deligne group ${\mathrm{WD}}_{F}$ and L-group ${}^{L}G$. First, thanks to the Local Langlands correspondence for ${\mathrm{GL}}_{n}(F)$ ([HT] and [Hen3]), the local converse theorem of Henniart or its sharp refinement implies that if $\phi:{\mathrm{WD}}_{F}\to{}^{L}G$ is any admissible homomorphism with associated Weil-Deligne parameter $(\varphi,N)$ (see Section 2.2), and $r:{}^{L}G\to{\mathrm{GL}}(V)$ is an irreducible representation of ${}^{L}G$, then the twisted gamma factors of $r\circ\phi$ determine the ${\mathrm{GL}}(V)$-conjugacy class of $r\circ\varphi$ (see Section 3). Suppose that this holds for any irreducible $r$, then, when $G$ is $F$-split, a classical result of Steinberg implies that the conjugacy class of $\varphi(\gamma)$ by the Langlands dual group $\widehat{G}$ is uniquely determined for any $\gamma\in\mathrm{W}_{F}$. Under the assumption that $\widehat{G}$ is acceptable, using in particular results of Yu on acceptability (Section 4) and of Silberger and Zink on the Langlands classification of parameters (Section 5), we prove in Theorem 5.3 that this property in turn implies that the equivalence class of $\varphi$ is uniquely determined. If $\phi$ is moreover assumed to be generic this determines the class of $\phi$ uniquely, this folklore fact having recently been proved rigorously in [DHKM] as we recall in Section 2.2. This leads to our main result Theorem 6.3, to which we refer for more complete statements. We also prove that the statement of Theorem 6.3 does not apply to the group ${\mathrm{SO}}_{6}(F)$ for some specific fields $F$, by studying unacceptable parameters in Section 7. Finally using ideas similar to those of our main theorem, we prove the variant which coincides with the known local converse theorems via the generic local Langlands correspondence in Theorem 8.2.
@@ Acknowledgement.
We thank an anonymous referee for encouraging us to meditate on the notion of local converse theorem. We warmly thank Guy Henniart, Wee-Teck Gan and Rob Kurinczuk for very useful correspondence. We especially thank Ahmed Moussaoui for answering many questions about this work, and Eyal Kaplan for very useful correspondence and detailed comments on an earlier draft, allowing significant improvement of its exposition. Finally and most importantly, we thank Dipendra Prasad for many useful answers and bringing the notion of acceptability to our attention, as well as for spotting a mistake in a previous version, the correction of which allowed to strengthen our results.
@@ 2.1. Langlands parameters and their local constants
and we say that two Weil-Deligne parameters are equivlent if they are in the same $\widehat{G}$-orbit. Let $\phi:{\mathrm{WD}}_{F}\to{}^{L}G$ be an admissible homomorphism, and denote by $d\phi$ the differential of $\phi_{|{\mathrm{SL}}_{2}({\mathbb{C}})}$ at $I_{2}$. As explained in [DHKM, Remark 6.4, (2)], one asociates to $\phi$ the Weil-Deligne parameter $(\varphi,N)$ where
@@ 3. A consequence of the sharp local converse theorem for ${\mathrm{GL}}_{n}$
Let $\sum_{i=1}^{t}\varphi_{i}\otimes{\mathrm{Sp}}(r_{i})$ be a semi-simple representation of ${\mathrm{WD}}_{F}$ with $\varphi_{i}$ trivial on $\mathrm{W}_{F}$. Set $\rho_{i}:=\mathrm{LLC}(\varphi_{i})$ and assume that the real numbers $e(\mathrm{St}_{r_{i}}(\rho_{i}))$ form a non increasing sequence. Then:
@@ 4. Acceptable reductive algebraic groups over ${\mathbb{C}}$
Moreover we mention that the references [Lar1], [Gri], [Lar2], [CG] and [Yu] provide a quite complete classification of acceptable groups. Now we provide a list of acceptable groups. Taking product of the groups in the list below provides other examples.
@@
The groups ${\mathrm{GL}}_{n}({\mathbb{C}})$, ${\mathrm{SL}}_{n}({\mathbb{C}})$, ${\mathrm{Sp}}_{2n}({\mathbb{C}})$, ${\mathrm{SO}}_{2n+1}({\mathbb{C}})$, $\mathrm{G}_{2}({\mathbb{C}})$ and ${\mathrm{SO}}_{4}({\mathbb{C}})$ are all acceptable.
On the other hand even orthogonal ${\mathrm{SO}}_{2n}({\mathbb{C}})$ for $n$ larger than $3$, are for example unacceptable.
@@ 6. A local converse theorem for Langlands parameters
    Suppose that for any irreducible representation $r$ of ${}^{L}G$, and any irreducible representation $\phi$ of ${\mathrm{WD}}_{K}$, one has
@@
    If $G$ is $F$-split, semi-simple and simply connected, it is sufficient to take $r$ above amongst the finite number of fundamental representations of ${}^{L}G$, and to twist by irreducible representations $\phi$ of ${\mathrm{WD}}_{K}$ of dimension at most $d$, where $d:=\lfloor N/2\rfloor$ for $N$ the dimension of any fundamental representation of highest dimension.
@@
This irreducible representation of ${}^{L}G$ (obtained by inflating that of its quotient ${\mathrm{GL}}_{n}({\mathbb{C}})\rtimes{\mathbb{Z}}/2{\mathbb{Z}}$) is enough for our modest purpose. Under our assumption, the restriction of $(\phi_{i})_{E}$ to of $\phi_{i}$ to ${\mathrm{WD}}_{E}$ is irreducible. From this it follows that $(r_{0}\circ\phi_{i})_{E}=(\phi_{i})_{E}\oplus\alpha\circ(\phi_{i})_{E}$, hence either $(\phi_{2})_{E}=(\phi_{1})_{E}$, or $(\phi_{2})_{E}=\alpha\circ(\phi_{1})_{E}$. We deduce that $\phi_{2}=\phi_{1}$, or $\phi_{2}=\alpha\circ\phi_{1}$ thanks to for example to [Pra, Proposition 7.10]. ∎
@@ 7. Unacceptable parameters
Suppose that $G$ is an $F$-split connected reductive group, and that there exists $\varphi_{1}:\mathrm{W}_{F}\to\widehat{G}$ which is unacceptable. Let $\varphi_{2}$ be locally conjugate but not globally conjugate to $\varphi_{1}$. Now take admissible homomorphisms $\phi_{1},\phi_{2}:\mathrm{W}_{F}\to{}^{L}G$ such that the semi-simple part of the Weil-Deligne parameter attached to $\phi_{1}$ is $\varphi_{1}$, and that of the Weil-Deligne parameter attached to $\phi_{2}$ is $\varphi_{2}$. Then for any irreducible representation $r$ of ${}^{L}G$, and any irreducible representation $\phi$ of ${\mathrm{WD}}_{F}$, one has
@@
Now we deduce from [Yu, Example 2.1] due to Weidner ([Wei]), the following statement. It is interesting to compare it with the remark ending the introduction of [YZ].
@@
Suppose that $14|q-$ 1. Then there exist non equivalent generic homomorphisms $\phi_{1},\phi_{2}:{\mathrm{WD}}_{F}\to{}^{L}{\mathrm{SO}}_{6}(F)$ such that for any irreducible representation $r$ of ${}^{L}G$, and any irreducible representation $\phi$ of ${\mathrm{WD}}_{F}$, one has
@@
Under the condition that $14|q-$ 1, and by class field theory, the group $(\frac{{\mathbb{Z}}}{4{\mathbb{Z}}})^{2}=\langle\gamma_{1},\ \gamma_{2}\rangle$ is a quotient of $\mathrm{W}_{F}$ because it is a quotient of $F^{\times}$. In particular let $\widetilde{\phi_{1}}:(\frac{{\mathbb{Z}}}{4{\mathbb{Z}}})^{2}\to{\mathrm{SL}}_{4}({\mathbb{C}})$ and $\widetilde{\phi_{2}}:(\frac{{\mathbb{Z}}}{4{\mathbb{Z}}})^{2}\to{\mathrm{SL}}_{4}({\mathbb{C}})$ the homomorphisms defined in [Yu, Example 2.1]: $\widetilde{\phi_{1}}(\gamma_{1})=\mathrm{diag}(1,1,i,-i)$, $\widetilde{\varphi_{1}}(\gamma_{2})=\mathrm{diag}(1,i,1,-i)$, and $\widetilde{\varphi_{2}}:=\overline{\widetilde{\varphi_{1}}}$. Set $\varphi_{i}$ to be the composition of $\widetilde{\phi_{i}}$ with the projection from ${\mathrm{SL}}_{4}({\mathbb{C}})$ to ${\mathrm{SL}}_{4}({\mathbb{C}})/\{\pm I_{4}\}\simeq{\mathrm{SO}}_{6}({\mathbb{C}})\subseteq{}^{L}{\mathrm{SO}}_{6}(F)$, and identify $\varphi_{i}$ with a semi-simple homomorphism $\mathrm{W}_{F}\to{}^{L}{\mathrm{SO}}_{6}(F)$ by inflation. One checks that the centralizer of both $\varphi_{1}$ and $\varphi_{2}$ are equal: seen in ${\mathrm{SL}}_{4}({\mathbb{C}})/\{\pm I_{4}\}$ they are equal to $D/\{\pm I_{4}\}$ where $D$ is the diagonal torus of ${\mathrm{SL}}_{4}({\mathbb{C}})$. In particular, according to [DHKM, Proposition 6.10], we can take a nilpotent element $N$ in the Lie algebra of ${\mathrm{SO}}_{6}({\mathbb{C}})$ such that $(\varphi_{1},N)$ and $(\varphi_{2},N)$ are Weil-Deligne parameters, which are moreover generic. We then take $\phi_{i}$ corresponding to $(\varphi_{i},N)$, and apply Proposition 7.2. ∎
An interesting question is to know if one can find such examples whenever $\widehat{G}$ is unacceptable, for some non Archimedean local field $F$, or even for all. This has to do with the classification of unacceptable homomorphisms.

### v5

@@ Abstract.
Let $F$ be a non Archimedean local field, and $G$ be the $F$-points of a connected quasi-split reductive group defined over $F$. In this paper we prove a converse theorem statement for generic Langlands parameters of $G$ when $G$ is $F$-split and the Langlands dual group of $G$ is acceptable, and propose a possible extension of it without the split assumption. We also prove that the statement does not apply to ${\mathrm{SO}}_{2n}(F)$ for certain choices of $F$, as soon as $n\geq 3$. Then we consider a variant which we prove for $G=\mathrm{G}_{2}(F)$ and all quasi-split classical groups. When $F$ has characteristic zero and assuming the validity of the Gross-Prasad and Rallis conjecture, this latter variant translates via the generic local Langlands correspondence of Jantzen and Liu, into the usual local converse theorems for classical groups expressed in terms of Shahidi’s gamma factors.
@@ 1. Introduction
We now briefly explain the main observations that are made in this paper. We denote by $G$ be the $F$-points of a connected quasi-split reductive group defined over $F$, Weil-Deligne group ${\mathrm{WD}}_{F}=\mathrm{W}_{F}\times{\mathrm{SL}}_{2}({\mathbb{C}})$ and L-group ${}^{L}G=G^{\vee}\ltimes\mathrm{W}_{F}$. First, thanks to the Local Langlands correspondence for ${\mathrm{GL}}_{n}(F)$ ([HT] and [Hen3]), the local converse theorem of Henniart or its sharp refinement implies that if $\phi:{\mathrm{WD}}_{F}\to{}^{L}G$ is any admissible homomorphism with associated Weil-Deligne parameter $(\varphi,N)$ (see Section 2.2), and $r:{}^{L}G\to{\mathrm{GL}}(V)$ is an irreducible representation of ${}^{L}G$, then the twisted gamma factors of $r\circ\phi$ determine the ${\mathrm{GL}}(V)$-conjugacy class of $r\circ\varphi$ (see Section 3). Suppose that this holds for any irreducible $r$, then, when $G$ is $F$-split, a classical result of Steinberg implies that the conjugacy class of $\varphi(\gamma)$ by the Langlands dual group $\widehat{G}$ is uniquely determined for any $\gamma\in\mathrm{W}_{F}$. Under the assumption that $\widehat{G}$ is acceptable, using in particular results of Yu on acceptability (Section 4) and of Silberger and Zink on the Langlands classification of parameters (Section 5), we prove in Theorem 5.3 that this property in turn implies that the equivalence class of $\varphi$ is uniquely determined. If $\phi$ is moreover assumed to be generic this determines the class of $\phi$ uniquely, this folklore fact having recently been proved rigorously in [DHKM] as we recall in Section 2.2. As a summary of this discussion, we state the key result of this paper. Say that $G$ is $\mathrm{W}_{F}$-acceptable if whenever we have parameters $\varphi,\varphi^{\prime}:\mathrm{W}_{F}\to\widehat{G}$ such that $r\circ\varphi$ is isomorphic to $r\circ\varphi^{\prime}$ for all algebraic representations $r$ of $\widehat{G}$, then $\varphi$ is isomorphic to $\varphi^{\prime}$. Say that $G$ is generically ${\mathrm{WD}}_{F}$-acceptable when the above property holds whenever the parameters $\varphi,\varphi^{\prime}:{\mathrm{WD}}_{F}\to\widehat{G}$ are generic.
Theorem.
Suppose that $\widehat{G}$ is acceptable, then $G$ is $\mathrm{W}_{F}$-acceptable, hence it is generically ${\mathrm{WD}}_{F}$-acceptable.
This leads to our main result Theorem 6.3, to which we refer for more complete statements. We also prove that the statement of Theorem 6.3 does not apply to the group ${\mathrm{SO}}_{2n}(F)$ for some specific fields $F$ as soon as $n\geq 3$, by studying unacceptable parameters in Section 7. An interesting question, which woud require a case by case analysis of unacceptable parameters but which can be addressed thanks to the methods of this paper, is whether local converse theorems for split groups only hold when the Langlands dual group is acceptable. Finally using ideas similar to those of our main theorem, we prove the variant which coincides with the known local converse theorems via the generic local Langlands correspondence in Theorem 8.2.
@@ Acknowledgement.
We thank the referee for many useful comments. We thank Moshe Adrian, Guy Henniart, Wee-Teck Gan, Rob Kurinczuk and Shaun Stevens for useful correspondence. We especially thank Ahmed Moussaoui for answering many questions about this work, and Eyal Kaplan for very useful correspondence and detailed comments on an earlier draft, allowing significant improvement of its exposition. Finally and most importantly, we thank Dipendra Prasad for many useful answers and bringing the notion of acceptability to our attention, as well as for spotting a mistake in a previous version, the correction of which allowed to strengthen our results.
@@ 2.1. Langlands parameters and their local constants
and we say that two Weil-Deligne parameters are equivalent if they are in the same $\widehat{G}$-orbit. Let $\phi:{\mathrm{WD}}_{F}\to{}^{L}G$ be an admissible homomorphism, and denote by $d\phi$ the differential of $\phi_{|{\mathrm{SL}}_{2}({\mathbb{C}})}$ at $I_{2}$. As explained in [DHKM, Remark 6.4, (2)], one asociates to $\phi$ the Weil-Deligne parameter $(\varphi,N)$ where
@@ 3. A consequence of the sharp local converse theorem for ${\mathrm{GL}}_{n}$
Let $\sum_{i=1}^{t}\varphi_{i}\otimes\mathrm{sp}(r_{i})$ be a semi-simple representation of ${\mathrm{WD}}_{F}$ with $\varphi_{i}$ trivial on ${\mathrm{SL}}_{2}({\mathbb{C}})$. Set $\rho_{i}:=\mathrm{LLC}(\varphi_{i})$ and assume that the real numbers $e(\mathrm{St}_{r_{i}}(\rho_{i}))$ form a non increasing sequence. Then:
@@ 4. Acceptable reductive algebraic groups over ${\mathbb{C}}$
Moreover we mention that the references [Lar1], [Gri], [Lar2], [CG] and [Yu] provide a quite complete classification of acceptable groups.
@@
A complex reductive group is admissible if and only if its derived subgroup is a product of groups in the following list: ${\mathrm{SL}}_{n}({\mathbb{C}})$, ${\mathrm{Sp}}_{2n}({\mathbb{C}})$, ${\mathrm{SO}}_{2n+1}({\mathbb{C}})$, $\mathrm{G}_{2}({\mathbb{C}})$ and ${\mathrm{SO}}_{4}({\mathbb{C}})$.
On the other hand even orthogonal ${\mathrm{SO}}_{2n}({\mathbb{C}})$ for $n$ larger than 333, are for example unacceptable.
@@ 6. A local converse theorem for Langlands parameters
    Suppose that for any irreducible representation $r$ of ${}^{L}G$, and any irreducible representation $\phi$ of ${\mathrm{WD}}_{F}$, one has
@@
    If $G$ is $F$-split, semi-simple and simply connected, it is sufficient to take $r$ above amongst the finite number of fundamental representations of ${}^{L}G$, and to twist by irreducible representations $\phi$ of ${\mathrm{WD}}_{F}$ of dimension at most $d$, where $d:=\lfloor N/2\rfloor$ for $N$ the dimension of any fundamental representation of highest dimension.
@@
This irreducible representation of ${}^{L}G$ (obtained by inflating that of its quotient ${\mathrm{GL}}_{n}({\mathbb{C}})\rtimes{\mathbb{Z}}/2{\mathbb{Z}}$) is enough for our modest purpose. Under our assumption, the restriction $(\phi_{i})_{E}$ of $\phi_{i}$ to ${\mathrm{WD}}_{E}$ is irreducible. From this it follows that $(r_{0}\circ\phi_{i})_{E}=(\phi_{i})_{E}\oplus\alpha\circ(\phi_{i})_{E}$, hence either $(\phi_{2})_{E}=(\phi_{1})_{E}$, or $(\phi_{2})_{E}=\alpha\circ(\phi_{1})_{E}$. We deduce that $\phi_{2}=\phi_{1}$, or $\phi_{2}=\alpha\circ\phi_{1}$ thanks to for example [Pra, Proposition 7.10]. ∎
@@ 7. Unacceptable parameters
Suppose that $G$ is an $F$-split connected reductive group, and that there exists $\varphi_{1}:\mathrm{W}_{F}\to\widehat{G}$ which is unacceptable. Let $\varphi_{2}$ be locally conjugate but not globally conjugate to $\varphi_{1}$. Now take admissible homomorphisms $\phi_{1},\phi_{2}:{\mathrm{WD}}_{F}\to{}^{L}G$ such that the semi-simple part of the Weil-Deligne parameter attached to $\phi_{i}$ is $\varphi_{i}$. Then for any irreducible representation $r$ of ${}^{L}G$, and any irreducible representation $\phi$ of ${\mathrm{WD}}_{F}$, one has
@@
Now we deduce from [Wei, Theorem 24], which we explicate further, the following statement. It is interesting to compare it with the remark ending the introduction of [YZ].
@@
Suppose that $4|q-1$ and that $n\geq 3$. Then there exist non equivalent generic homomorphisms $\phi_{1},\phi_{2}:{\mathrm{WD}}_{F}\to{}^{L}{\mathrm{SO}}_{2n}(F)$ such that for any irreducible representation $r$ of ${}^{L}G$, and any irreducible representation $\phi$ of ${\mathrm{WD}}_{F}$, one has
@@
Under the condition that $4|q-1$, and by class field theory, the group $(\frac{{\mathbb{Z}}}{4{\mathbb{Z}}})^{2}$ is a quotient of $\mathrm{W}_{F}$ because it is a quotient of $F^{\times}$. In particular [Wei, Theorem 24] provides locally but non globally conjugate parameters $\varphi_{i}:\mathrm{W}_{F}\to{}^{L}{\mathrm{SO}}_{2n}(F)$, $i=1,2$, by inflation. Let us slightly rephrase and explicate Weidner’s example. We denote by $w_{n}$ the antidiagonal matrix in ${\mathrm{GL}}_{n}({\mathbb{C}})$ with ones on the second diagonal. We realize ${}^{L}{\mathrm{SO}}_{2n}(F)={\mathrm{SO}}_{2n}({\mathbb{C}})$ as the group
${\mathrm{SO}}_{2n}({\mathbb{C}}):=\{M\in{\mathrm{GL}}_{2n}({\mathbb{C}}),\ w_{2n}M^{T}w_{2n}^{-1}=M^{-1}\}.$
We define $\varphi_{i}:\mathrm{W}_{F}\to{\mathrm{SO}}_{2n}({\mathbb{C}})$ as follows. We write
$F^{\times}=\varpi_{F}^{{\mathbb{Z}}}\times\langle u\rangle\times(1+\varpi_{F}O_{F})$
where $\varpi_{F}$ is a uniformizer of $F$, $O_{F}$ is the ring of integers of $F$, and $u$ has order $q-1$. Then
$\varphi_{i}:\mathrm{W}_{F}\to F^{\times}\overset{\lambda_{i}}{\to}{\mathrm{SO}}_{2n}({\mathbb{C}}),$
where $\lambda_{1}$ is trivial on $1+\varpi_{F}O_{F}$,
$\lambda_{1}(\varpi_{F})=\mathrm{diag}(iI_{n-1},I_{2},-iI_{n-1})$
and
$\lambda_{1}(u)=\mathrm{diag}(1,iI_{n-1},-iI_{n-1},1),$
whereas
$\lambda_{2}:=\mathrm{diag}(I_{n-1},w_{2},I_{n-1})\lambda_{1}\mathrm{diag}(I_{n-1},w_{2},I_{n-1})^{-1}.$
Since
$\lambda_{1}(\varpi_{F}^{a}u^{b})=\mathrm{diag}(i^{a},i^{a+b}I_{n-2},i^{b},(-i)^{b},(-i)^{a+b}I_{n-2},(-i)^{a})$
and
$\lambda_{2}(\varpi_{F}^{a}u^{b})=\mathrm{diag}(i^{a},i^{a+b}I_{n-2},(-i)^{b},i^{b},(-i)^{a+b}I_{n-2},(-i)^{a}),$
we see that $\varphi_{1}$ and $\varphi_{2}$ are locally conjugate in ${\mathrm{SO}}(2n,{\mathbb{C}})$ by examining the parity of $a$ and $b$:
  • when $b$ is even then $\lambda_{1}(\varpi_{F}^{a}u^{b})=\lambda_{2}(\varpi_{F}^{a}u^{b})$,
[…]
  • when $a$ and $b$ are odd then $\lambda_{1}(\varpi_{F}^{a}u^{b})$ and $\lambda_{2}(\varpi_{F}^{a}u^{b})$ are conjugate by the permutation matrix corresponding to the double transposition $(n\ n+1)\circ(n-1\ n+2)$.
On the other hand, if they were globally conjugate, they would be conjugate by an element in the centralizer of $\lambda_{1}(\varpi_{F})$, which is the set of matrices
$\{\mathrm{diag}(A,B,w_{n-1}A^{-T}w_{n-1}),\ A\in{\mathrm{GL}}_{n-1}({\mathbb{C}}),\ B\in{\mathrm{SO}}_{2}({\mathbb{C}})\}.$
This is impossible since $\mathrm{diag}(i,-i)$ is not ${\mathrm{SO}}_{2}({\mathbb{C}})$-conjugate to $\mathrm{diag}(-i,i)$, as ${\mathrm{SO}}_{2}({\mathbb{C}})$ consists of diagonal matrices with determinant one. Moreover the centralizers of $\varphi_{1}$ and $\varphi_{2}$ are both equal to the set
$\{\mathrm{diag}(z,A,B,w_{n-2}A^{-T}w_{n-2},z^{-1}),\ A\in{\mathrm{GL}}_{n-2}({\mathbb{C}}),\ B\in{\mathrm{SO}}_{2}({\mathbb{C}}),\ z\in{\mathbb{C}}^{\times}\}.$
In particular, according to [DHKM, Proposition 6.10], we can take a nilpotent element $N$ in the Lie algebra of ${\mathrm{SO}}_{2n}({\mathbb{C}})$ such that $(\varphi_{1},N)$ and $(\varphi_{2},N)$ are Weil-Deligne parameters, which are moreover generic. We then take $\phi_{i}$ corresponding to $(\varphi_{i},N)$, and apply Proposition 7.2. ∎
Remark 7.4.
The monodromy operator $N$ in the proof above must be equal to zero since the eigenvalues of the adjoint action of $(\frac{{\mathbb{Z}}}{4{\mathbb{Z}}})^{2}$ on the Lie algebra of $\widehat{G}$ are of absolute value one, whereas the image of the Frobenius element in the group $(\frac{{\mathbb{Z}}}{4{\mathbb{Z}}})^{2}$ would act with eigenvalue $q^{-1}$ on $N$ if $N$ was nonzero.
Remark 7.5.
An interesting question is to know if one can find such examples whenever $\widehat{G}$ is unacceptable, for some non Archimedean local field $F$, or even for all. This has to do with the classification of unacceptable homomorphisms. Namely, the question is: is it possible to find for any unacceptable $\widehat{G}$, an unacceptable parameter $\varphi:\Gamma\to\widehat{G}$ such that $\Gamma$ is a quotient of $\mathrm{W}_{F}$?